OpenStudy (anonymous):

Find the diagonal of a cube if its side equals 5.

OpenStudy (anonymous):

 using Pythagoras theorem will help you find the diagonal \[a^2 + b^2 = c^2\] so if we look at the triangle we have two legs both equal to 5 and one hypot = c so plug in the given information and we got \[5^2 + 5^2 = c^2\] can you take it from here?